OpenStudy (anonymous):

Please I need help ASAP! Got all my math done during spring break but need help with these two questions!! Thanks a bunch and will medal!! Questions to follow

OpenStudy (anonymous):

Sorry put in wrong Question this is correct ones A researcher tagged 40 giant turtles in 2008. In 2009, she returned and captured 50 turtles, 16 of which were tagged. In 2010, she returned again and captured 30 turtles, 8 of which were tagged. a. Use a proportion to estimate the turtle population for 2009. b. Use a proportion to estimate the turtle population for 2010.

OpenStudy (anonymous):

a) Describe the overlap of the two sets of data. b) Calculate the mean of each data set. c) Calculate the mean absolute deviation (MAD) of each data set. d) Which set is more variable? How do you know?

OpenStudy (anonymous):

I think on second question b.) is 71.8, 72 as mean for both data sets

OpenStudy (goformit100):

@stuckinarut you r correct. You are Genius... you have Selected the correct option.

OpenStudy (anonymous):

You need to figure out what proportion you are taking and how that proportion will help you estimate the turtle population. Would you guess there are more turtles if a) all the turtles caught in 2009 were tagged or b) none of them were? Why?

OpenStudy (anonymous):

Not stupid at doing the math just dont understand how to correctly get the problem started on either of these but hey thanks @goformit100

OpenStudy (goformit100):

@Dharmaputra_993 Sir plz Hel Him.

OpenStudy (goformit100):

My Pleasure @stuckinarut

OpenStudy (anonymous):

@goformit100 Hel him?

OpenStudy (anonymous):

I thought 34:50 for 2009

OpenStudy (anonymous):

That could work. What is that a proportion of?

OpenStudy (anonymous):

untagged turtles

OpenStudy (anonymous):

Well, let's say x is the actual turtle population. 16 out of 50 turtles were tagged, and there were 40 tagged turtles originally. Then we could assume that the following equation holds: \[\frac{16}{50} = \frac{40}{x}\]

OpenStudy (anonymous):

because the proportion of the tagged turtles should be about the same, whether we're looking at only 50 turtles or at the whole population

OpenStudy (anonymous):

Compared to what? Proportions always have two parts, right. Then you have to compare that to another proportion comparing the same information.

OpenStudy (anonymous):

Ah gotcha!!

OpenStudy (anonymous):

You can use either 16 or 34, but you have to be careful what each of these numbers represents.

OpenStudy (anonymous):

Thats what kinda got me too

OpenStudy (anonymous):

well, if you used 34, you would have to write it like this:\[\frac{34}{50} = \frac{x-40}{x}\]

OpenStudy (anonymous):

but of course it's easier to solve the equation if you do it the "16-way"

OpenStudy (anonymous):

lol yea 125

OpenStudy (anonymous):

I think that's correct, got the same result.

OpenStudy (anonymous):

yea!!

OpenStudy (anonymous):

What about problem two? How do I find part b?

OpenStudy (anonymous):

Oh yea and on part b of problem 2 I believe that is 150 turtles

OpenStudy (anonymous):

Seems right. I'm not quite sure how to interpretate this data. On the "Soil A"-side, it says Key: 9|6 means 69, whereas on the "Soil-B"-side it says Key: 5|8 means 58. Doesn't really make a lot of sense. Do you think the first one is a typing error?

OpenStudy (anonymous):

no on the graph you will read the left side backwards so the middle number is first on them as well does that make sense

OpenStudy (anonymous):

ahh, ok. Now I get it

OpenStudy (anonymous):

Well, for the mean value of the left-hand side for example, you just add up all the data you've got, and then divide by the number of terms you added up

OpenStudy (anonymous):

got that part I believe its the part b that I dont know how to do

OpenStudy (anonymous):

or should say the steps to do it

OpenStudy (anonymous):

Do you think you would get it if I would describe it using mathematical formulas rather than words? I can't help it, I'm a mathematician

OpenStudy (anonymous):

yes that would be fine!

OpenStudy (anonymous):

let's say your data consists of n different numbers, I'll refer to them as\[x_1, x_2,...,x_i,...,x_n\]

OpenStudy (anonymous):

and let X be the mean value

OpenStudy (anonymous):

then the mean absolute deviation is calculated using the following formula: \[\frac{1}{n}\sum_{i=1}^{n} |X-x_i|\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Got it thank you so much for your help means ALOT!! Very helpful I really do appreciate your time

OpenStudy (anonymous):

you're welcome